In a Math test, the mean score is 42 and the standard deviation is 3. Assuming normality, what is the probability that the score picked will lie above score 50? *

1
Expert's answer
2022-04-06T09:40:32-0400

We have a normal distribution, "\mu=42, \sigma=3."

Let's convert it to the standard normal distribution,

"z=\cfrac{x-\mu}{\sigma}=\cfrac{50-42}{3}=2.67;"


"P(X>50)=P(Z>2.67)=\\=1-P(Z<2.67)="

"=1-0.9962=0.0038" ​(from z-table).